A bullet on penetrating 2 successive wooden planks of unqual thickness loses its velocity by 200meter per second in each case. If the initial velocity is 1000meter per second calculate the ratio of thickness of the Planks.

Grade:11

2 Answers

Eshan
5 years ago
Loss in kinetic energy of bulletthrough a block\proptothickness of the block.

Hence for the first block

\dfrac{1}{2}m(1000^2-200^2)\propto t_1

for the second block,

\dfrac{1}{2}m(200^2-0^2)\propto t_2

\implies \dfrac{t_1}{t_2}=\dfrac{1000^2-200^2}{200^2}=24
Ankan Pal
4 years ago
Initial velocity-1000 m/S
After penetrating 1st block final velocity -(1000-200)=800 m/S
For 2nd block initial velocity-800m/S
Final velocity-600 m/S
Now,
V1^2=u1^2-2as1
2as1=u^2-v^2.       _(1)
as1=18000
V2^2=u2^2-2as2
2as2=u2^2-v2^2
as2=14000.          _(2)
Now (1)÷(2)= as1÷as2=18000÷14000
S1:s2= 9:7
Hence the ratio of the thickness of the planks are 9:7